# Fine-tuning Diffusion Policies with Backpropagation Through Diffusion Timesteps

### v2

@@ Abstract
Diffusion policies, widely adopted in decision-making scenarios such as robotics, gaming and autonomous driving, are capable of learning diverse skills from demonstration data due to their high representation power. However, the sub-optimal and limited coverage of demonstration data could lead to diffusion policies that generate sub-optimal trajectories and even catastrophic failures. While reinforcement learning (RL)-based fine-tuning has emerged as a promising solution to address these limitations, existing approaches struggle to effectively adapt Proximal Policy Optimization (PPO) to diffusion models. This challenge stems from the computational intractability of action likelihood estimation during the denoising process, which leads to complicated optimization objectives. In our experiments starting from randomly initialized policies, we find that online tuning of Diffusion Policies demonstrates much lower sample efficiency compared to directly applying PPO on MLP policies (MLP+PPO). To address these challenges, we introduce NCDPO, a novel framework that reformulates Diffusion Policy as a noise-conditioned deterministic policy. By treating each denoising step as a differentiable transformation conditioned on pre-sampled noise, NCDPO enables tractable likelihood evaluation and gradient backpropagation through all diffusion timesteps. Our experiments demonstrate that NCDPO achieves sample efficiency comparable to MLP+PPO when training from scratch, outperforming existing methods in both sample efficiency and final performance across diverse benchmarks, including continuous robot control and multi-agent game scenarios. Furthermore, our experimental results show that our method is robust to the number denoising timesteps.
@@ 1 Introduction
Currently, the most effective approach, DPPO (Diffusion Policy Policy Optimization[20] employs Policy Gradient (PG) approaches to enhance the performance of pre-trained Diffusion Policy in continuous control tasks. By treating the denoising process of Diffusion Policy as a low-level Markov Decision Process, DPPO optimizes the Gaussian likelihood of all denoising steps. However, through our extensive experiments, we find fine-tuning Diffusion Policies with RL faces a challenge of sample efficiency. Specifically, in our RL experiments starting from randomly initialized policies, we find that training Diffusion Policy with DPPO could lead to worse sample efficiency and final performance than training an MLP policy with standard RL. We hypothesize that the training efficiency gap occurs since DPPO uses a much larger action space for RL training, which impedes the sample efficiency of RL training. Therefore, a question becomes particularly important: Can we design a more effective fine-tuning approach for Diffusion Policy that avoids enlarging the action space during RL training?
In this work, we present Noise-Conditioned Diffusion Policy Optimization (NCDPO), a sample-efficient RL algorithm for fine-tuning Diffusion Policies. NCDPO formulates the denoising process of Diffusion Policy as a noise-conditioned inference process, ensuring the RL objective only contain the likelihood of the interactive actions, i.e. actions generated by Diffusion Policy to interact with the environment. In the policy update phase, the gradients with respect to the policy parameters are computed with Backpropagation through Diffusion Timesteps (BPDT). When performing RL training on randomly initialized policies, we show that training Diffusion Policy with NCDPO achieves comparable sample efficiency with training an MLP policy with RL.
In summary, our main contribution is NCDPO, a novel framework which is applicable to both continuous and discrete environments, to fine-tune Diffusion Policies, by formulating denoising steps as deterministic generation process and apply PPO. We also evaluate NCDPO on a set of environments, ranging from continuous robot control and multi-agent coordination tasks. We demonstrate that NCDPO obtains higher sample efficiency and stronger final performance than baseline methods across all evaluated environments. Finally, our ablation study reveals that that NCDPO is robust to the number of diffusion timesteps and remains highly sample efficient when the number of diffusion timesteps is large.
@@ 2 Related Work
Diffusion-based generative models have demonstrated remarkable effectiveness in the domains of visual content generation [23, 26, 19].One central capability of Diffusion Models is the denoising process that iteratively refines sampled noises into clean datapoints. [9, 24, 25]. Beyond their success in content generation, diffusion models have increasingly been adapted for decision-making tasks across a range of domains, including robotics [5, 21, 2, 33, 27, 15], autonomous driving [13, 30], and gaming [15, 34]. In robotics, most existing work trains Diffusion Policies through imitation learning. For instance, Reuss et al. [21] predict future action chunks using goal-conditioned imitation learning, while [33] integrate Diffusion Policies with compact 3D representations extracted from point clouds. To further enhance the quality of generated behaviors, return signal or goal conditioning is applied to encourage the generation of high-value actions [10, 1, 12].
@@
Recent works have aimed to enhance learned Diffusion Policy through fine-tuning with Reinforcement Learning approaches. A line of work has been focusing on integrating Diffusion Policies with Q-learning using offline data [4, 11, 28, 7, 22, 35, 18]. In addition to offline reinforcement learning, recent advancements have explored fine-tuning Diffusion Policies with online RL algorithms, for example, aligning the score function with the action gradient [31], or employing the diffusion model as a policy extraction mechanism within implicit Q-learning [8]. Most recently, [20] formulates the denoising process of Diffusion Policy as a "Diffusion MDP", enabling the application of RL algorithms to optimize all denoising steps with online feedback. In this work, we investigate an alternative representation for the denoising process that enables sample efficient fine-tuning of Diffusion Policy.
@@ 3 Preliminary
where $\sigma$ is a fixed noise schedule for action generation, $\beta$ denotes the forward process variances and is held as constant, and $\theta$ is the parameter of Diffusion Policy. To avoid ambiguity, we use interactive actions to denote the action $a^{0}$ that is used for interacting with the environment and latent actions to denote actions $a^{1},\cdots,a^{K}$ that are generated during the denoising process. For more training details on diffusion models, please refer to [9].
@@
[…]
@@ 4 Sample Efficiency Challenge for Diffusion Policy Fine-Tuning
In this section, we aim to investigate the sample efficiency of fine-tuning Diffusion Policy with RL. Specifically, we compare the sample efficiency of training Diffusion Policy using DPPO and training an MLP policy using standard PPO. Since our study only focuses on the RL process, both the MLP policy and Diffusion Policy are randomly initialized before RL training without performing any additional behavior cloning. For conciseness, we denote training Diffusion Policy with DPPO as DP+DPPO and training an MLP policy with PPO as MLP+PPO.
Our investigations are carried out on two OpenAI Gym locomotion tasks, Walker2D and Halfcheetah. The training curves of MLP+PPO and DP+DPPO are shown in Fig. 1. Although Diffusion Policy has more powerful representation power than MLP policy [5], our results here surprisingly show that DP+DPPO is less sample efficient than MLP+PPO and could only achieves sub-optimal performance.
Why does this efficiency gap occur? We hypothesize that the underlying reason is that, by employing a two-level MDP formulation, DPPO actually significantly lengthens the MDP horizons in RL training to contain both the interactive actions and latent actions. This lengthened MDP horizon then results in difficulty in proper credit assignment. This insight raises a critical question: Can we design an alternative RL algorithm for Diffusion Policy fine-tuning that avoids enlarging the action space?
@@ 5 Noise-Conditioned Diffusion Policy Optimization
As discussed in Sec. 4, fine-tuning Diffusion Policy with a two-level MDP formulation could lead to sub-optimal sample efficiency and final performance. In this section, we present a novel sample-efficient RL training method for Diffusion Policy, Noise-Conditioned Diffusion Policy Optimization (NCDPO). In Sec. 5.1, we show that NCDPO formulates the denoising process of Diffusion Policy as a noise-conditioned inference process. In Sec. 5.2, we show that NCDPO ensures PPO training operates on the same action space as the environment, without relying on optimizing action likelihood of latent actions.
@@ 5.1 Denoising Process as a Noise-Conditioned Inference Process
We decouple the stochastic and deterministic components of the denoising process. The stochastic component encompasses all the random noises sampled during the denoising process. The deterministic component further operates on these sampled noises with the model $\mu_{\theta}$.
Formally, Eq. 4 can be equivalently represented as,
$a^{k-1}=\mu_{\theta}(a^{k},k,s)+\sigma_{k}\cdot z^{k}\;\;\text{ where\ }z^{k}\sim\mathcal{N}(0,I)$ (5)
A straitforward indication of Eq. 5 is that, in each denoising step, the only stochastic component is the Gaussian noise $z^{k}$, while the computation of $\mu_{\theta}(a^{k},k,s)$ and addition between $\mu_{\theta}(a^{k},k,s)$ and $\sigma_{k}\cdot z^{k}$ are both deterministic. Therefore, the whole denoising process can be split into a noise sampling phase and a deterministic inference phase.
In the noise sampling phase, we generate a sequence of standard Gaussian noises $z^{1},\cdots,z^{K}$,
$z^{k}\sim\mathcal{N}(0,I)\;\;\text{ for }k=K,K-1,\dots,1$ (6)
$a^{K}\sim\mathcal{N}(0,I)$ (7)
In the deterministic inference phase, given noises $z^{1},\cdots,z^{K}$, $\mu_{\theta}$ is be used to compute the latent actions $a^{k}$ one by one. For $k=K,K-1,\dots,1$, $a^{k-1}$ is a linear combination of $\mu_{\theta}(a^{k},k,s)$ and $z^{k}$,
$a^{k-1}=\mu_{\theta}(a^{k},k,s)+\sigma_{k}\cdot z^{k}$ (8)
Consequently, the generated action $a^{0}$ can be computed by recursively applying Eq. 8,
@@
MDP with Noise-augmented States.
As derived in Eq. 8 and Eq. 9, the denoising process can be partitioned into a noise sampling phase and a policy inference phase. We can incorporate the sampled noises into the MDP as part of the environment state. Formally, for the original environment MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P_{0},P,R,\gamma\rangle$ we introduce MDP with noise-augmented states $\mathcal{M}_{noise}=\langle\mathcal{S}_{noise},\mathcal{A},P_{0},P_{noise},R,\gamma\rangle$. In $\mathcal{M}_{noise}$, each state $s_{noise}$ consists of a environment state $s\in\mathcal{S}$ and Gaussian noises $a^{K},z^{1},\cdots,z^{K}$. $\mathcal{M}_{noise}$ shares the same action space and reward function as the original MDP $\mathcal{M}$. In each decision-making step, the environment state transits to a new one, and the noises are all re-sampled.
@@
Figure 2: DPPO adopts a two-layer MDP design by combining the environment state with latent actions to form augmented states. In contrast, in each step, NCDPO first samples a group of random noises and computes the action based on the noises, resulting in a deterministic generation process (Eq. 9). Blue arrows in the figure indicate MDP transitions.
[.45]Refer to caption [.5]Refer to caption
Denoising Process as a Noise-Conditioned Policy.
@@
Given a noise-augmented state $s_{noise}=(s,a^{K},z^{1:K})$, the deterministic inference phase of the denoising process can be represented as a Noise-Conditioned Policy $\pi^{NC}_{\theta}$ that generates the action $a^{0}$ using Eq. 9
@@
Note that this noise-conditoned policy is a deterministic policy and could not be directly trained with PPO since the policy loss (Eq. 2) relies on a stochastic policy. Therefore, we introduce an additional operation to transform this deterministic policy into a stochastic one. Specifically, for continuous action space, we sample the final action $a^{0}$ near $f(a^{K},z^{1:K})$ from a Gaussian distribution with a learnable standard variation $\sigma_{act}$,
$\pi^{NC}_{\theta}(\cdot|z^{1:K},a^{K},s)=\mathcal{N}(f_{\theta}(s,a^{K},z^{1:K}),\sigma_{act}^{2})$ (10)
For discrete action space, we use softmax to sample the final action by treating $f_{\theta}(s,a^{K},z^{1:K})$ as logits,
$\pi^{NC}_{\theta}(a^{0}=i|z^{1:K},a^{K},s)\propto{\exp{(f_{\theta}(s,a^{K},z^{1:K})_{i}/T)}}$ (11)
where $i$ is the action index. $T$ is the temperature that allows the policy network to produce sharper action distributions.
5.2 Finetuning Noised-Conditioned Policy with PPO
Under the formulation of NCDPO, at each denoising timestep $t$, Gaussian noise $z^{t}$ is first sampled, and the action is then generated via Eq. 10 or Eq. 11. This way, we can apply PPO objective in Eq.12 , which utilizes a clipped objective to regularize updated policy from original policy, to optimize interactive action probabilities:
@@ 5.2 Finetuning Noised-Conditioned Policy with PPO
$\begin{split}L(\theta|\theta_{k})=\mathbb{E}_{\begin{subarray}{c}a_{t}^{0}\sim\pi^{NC}_{\theta_{k}}(z_{t}^{1:K},a_{t}^{K},s_{t})\\ \end{subarray}}&\Biggl{[}\sum_{t}\min\Biggl{(}\frac{\pi^{NC}_{\theta}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})}{\pi^{NC}_{\theta_{k}}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})}A^{\pi^{NC}_{\theta_{k}}}(a_{t}^{0}|s_{t}),\\[2.84526pt] &\text{clip}\Bigl{(}\frac{\pi^{NC}_{\theta}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})}{\pi^{NC}_{\theta_{k}}(a_{t}^{0}|z_{t}^{1:K},a_{t}^{K},s_{t})},1-\epsilon,1+\epsilon\Bigr{)}A^{\pi^{NC}_{\theta_{k}}}(a_{t}^{0}|s_{t})\Biggr{)}\Biggr{]}\end{split}$ (12)
@@
As illustrated in Figure 16, in policy rollout process, each step begins by sampling a sequence of noises, which are then used by the Diffusion Policy to generate the corresponding action. These sampled noises are stored in the buffer. During training phase, the stored noises are reused to recompute the actions, enabling gradient backpropagation through the entire denoising process. This allows PPO to directly update all denoising steps of the diffusion policy.

### v4

@@ Abstract
Diffusion policies, widely adopted in decision-making scenarios such as robotics, gaming and autonomous driving, are capable of learning diverse skills from demonstration data due to their high representation power. However, the sub-optimal and limited coverage of demonstration data could lead to diffusion policies that generate sub-optimal trajectories and even catastrophic failures. While reinforcement learning (RL)-based fine-tuning has emerged as a promising solution to address these limitations, existing approaches struggle to effectively adapt Proximal Policy Optimization (PPO) to diffusion models. This challenge stems from the computational intractability of action likelihood estimation during the denoising process, which leads to complicated optimization objectives. In our experiments starting from randomly initialized policies, we find that online tuning of Diffusion Policies demonstrates much lower sample efficiency compared to directly applying PPO on MLP policies (MLP+PPO). To address these challenges, we introduce NCDPO, a novel framework that reformulates Diffusion Policy as a noise‐conditioned deterministic policy. By treating each denoising step as a differentiable transformation conditioned on pre‐sampled noise, NCDPO enables tractable likelihood evaluation and gradient backpropagation through all diffusion timesteps. This formulation enables direct optimization over the final denoised interactive actions without increasing MDP lengths. Our experiments demonstrate that NCDPO achieves sample efficiency comparable to MLP+PPO when training from scratch, outperforming existing methods in both sample efficiency and final performance across diverse benchmarks, including continuous robot control (with both state and vision inputs) and discrete multi-agent coordination tasks. Furthermore, our experimental results show that our method is robust to the number denoising timesteps.
@@ 1 Introduction
Currently, the most effective approach, DPPO (Diffusion Policy Policy Optimization[24] employs Policy Gradient (PG) approaches to enhance the performance of pre-trained Diffusion Policy in continuous control tasks. By treating the denoising process of Diffusion Policy as a low-level Markov Decision Process, DPPO optimizes the Gaussian likelihood of all denoising steps. However, through our extensive experiments, we find fine-tuning Diffusion Policies with RL faces a challenge of sample efficiency. Specifically, in our RL experiments starting from randomly initialized policies, we find that training Diffusion Policy with DPPO could lead to worse sample efficiency and final performance than training an MLP policy with standard RL. We hypothesize that the training efficiency gap occurs because DPPO uses a much longer MDP horizon for RL training, which impedes the sample efficiency of RL training. Therefore, a question becomes particularly important: Can we design a more effective fine-tuning approach for Diffusion Policy that avoids lengthening the MDP horizon during RL training?
In this paper, we present Noise-Conditioned Diffusion Policy Optimization (NCDPO), a sample-efficienct reinforcement learning algorithm for fine-tuning Diffusion Policies without increasing the MDP horizon. Specifically, we first interpret the denoising process as a noise-conditioned inference procedure: we pre-sample the complete noise sequence for all diffusion steps and treat it as fixed, making the denoising trajectory deterministic and allowing exact backpropagation. This enables us to reformulate the RL objective to apply PPO only to interactive actions—the fully denoised actions used for environment interaction—and apply Backpropagation through Diffusion Timesteps (BPDT) to propagate the PPO gradients backward through the entire denoising chain.
We also conduct extensive experiments across diverse domains, including continuous robot control with both state and vision inputs, as well as discrete multi-agent coordination. In all settings, NCDPO consistently surpasses baselines in both sample efficiency and final performance. On vision-based manipulation tasks, it achieves up to $4\times$ higher sample efficiency than DPPO, highlighting its promise for real-world applications. Ablation studies further demonstrate that NCDPO is robust to the choice of diffusion timesteps. We attribute these improvements to the shorter effective MDP length and the more accurate gradient estimation enabled by BPDT
@@ 2 Related Work
Diffusion-based generative models have demonstrated remarkable effectiveness in the domains of visual content generation [27, 30, 23].One central capability of Diffusion Models is the denoising process that iteratively refines sampled noises into clean datapoints [13, 28, 29]. Beyond their success in content generation, diffusion models have increasingly been adapted for decision-making tasks across a range of domains, including robotics [5, 25, 2, 36, 31, 19], autonomous driving [17, 33], and gaming [19, 37]. In robotics, most existing work trains Diffusion Policies through imitation learning. For instance, Reuss et al. [25] predict future action chunks using goal-conditioned imitation learning, while [36] integrate Diffusion Policies with compact 3D representations extracted from point clouds. To further enhance the quality of generated behaviors, return signal or goal conditioning is applied to encourage the generation of high-value actions [14, 1, 16].
@@
Recent works have aimed to enhance learned Diffusion Policy through fine-tuning with Reinforcement Learning approaches. A line of work has been focusing on integrating Diffusion Policies with Q-learning using offline data [4, 15, 32, 10, 26, 38, 22, 9, 6]. In addition to offline reinforcement learning, recent advancements have explored fine-tuning Diffusion Policies with online RL algorithms, for example, aligning the score function with the action gradient [34], or employing the diffusion model as a policy extraction mechanism within implicit Q-learning [12]. Most recently, Ren et al. [24] formulates the denoising process of Diffusion Policy as a "Diffusion MDP", enabling the application of RL algorithms to optimize all denoising steps with online feedback. In this work, we investigate an alternative representation for the denoising process that enables sample efficient fine-tuning of Diffusion Policy.
@@ 3 Preliminary
where $\sigma$ is a fixed noise schedule for action generation, $\beta$ denotes the forward process variances and is held as constant, and $\theta$ is the parameter of Diffusion Policy. To avoid ambiguity, we use interactive actions to denote the action $a^{0}$ that is used for interacting with the environment and latent actions to denote actions $a^{1},\cdots,a^{K}$ that are generated during the denoising process. For more training details on diffusion models, please refer to Ho et al. [13].
@@
$\pi_{\theta}(a_{t}^{0}|s_{t})=\int_{a^{1}_{t},\cdots,a^{K}_{t}}\mathbb{P}[a^{0}_{t},\cdots,a^{K}_{t}|s_{t},\pi_{\theta}]\cdot da^{1}_{t}\cdots da^{K}_{t}$
@@ 4 Sample Efficiency Challenge for Diffusion Policy Fine-Tuning
In this section, we study the sample efficiency of fine-tuning Diffusion Policies with reinforcement learning using one existing method. Concretely, we compare (i) training a diffusion-based policy with DPPO (denoted DP+DPPO) and (ii) training a standard MLP policy with PPO (denoted MLP+PPO). To isolate the RL process, both policies are randomly initialized before RL and no behavior cloning is used.
Experiments on two OpenAI Gym locomotion tasks (Walker2D and HalfCheetah) show that, despite the stronger representational power of Diffusion Policies [5], DP+DPPO is substantially less sample efficient than MLP+PPO and typically converges to a lower final performance (Fig. 1).
We hypothesize the reason is that, by employing a two-level MDP formulation, DPPO significantly lengthens the MDP horizons in RL training to contain both the interactive actions and latent actions. This lengthened MDP horizon then results in difficulty in proper credit assignment. This observation prompts us to design a different method that could fine-tune Diffusion Policies without lengthening the MDP horizon.
@@ 5 Noise-Conditioned Diffusion Policy Optimization
As motivated in Sec. 4, treating the diffusion denoising process as part of the MDP can harm sample efficiency. In this section, we present a novel sample-efficient RL training method for Diffusion Policy, Noise-Conditioned Diffusion Policy Optimization (NCDPO). NCDPO introduces backpropagation through diffusion timesteps by treating the diffusion denoising process as deterministic inference conditioned on pre-sampled noise, and applies PPO in the interactive action space. In Sec. 5.1 we formalize the noise-conditioned generative view. In Sec. 5.2 we show how PPO can be directly applied to the interactive actions instead of latent actions.
@@ 5.1 Denoising Process as a Noise-Conditioned Inference Process
Let $K$ be the number of denoising steps. A single denoising step can be written as
$a^{k-1}=\mu_{\theta}(a^{k},k,s)+\sigma_{k}\,z^{k},\qquad z^{k}\sim\mathcal{N}(0,I),$ (5)
for $k=K,K-1,\dots,1$, with $a^{K}\sim\mathcal{N}(0,I)$. The only stochastic terms are the Gaussian noises $a^{K}$ and $z^{1:K}$, and the map $\mu_{\theta}$ is deterministic given its inputs. Thus the sampling procedure naturally splits into a noise sampling phase and a deterministic inference phase.
In the noise sampling phase we draw
$a^{K}\sim\mathcal{N}(0,I),\qquad z^{k}\sim\mathcal{N}(0,I)\ \text{ for }k=1,\dots,K.$ (6)
In the deterministic inference phase we recursively apply Eq. (5) to obtain $a^{K-1},\dots,a^{0}$. Therefore the final action can be written as a deterministic function of the environment state and the sampled noises:
@@
Figure 2: DPPO adopts a two-layer MDP design by combining the environment state with latent actions to form augmented states. In contrast, in each step, NCDPO only applies PPO loss to the final interactive actions, resulting in shorter MDP lengths. (Eq. 10). Green arrows in the figure indicate gradient flow.
[.45]Refer to caption [.45]Refer to caption
MDP with Noise-augmented States.
As derived in Eq. 7, the denoising process can be divided into a noise sampling phase and a policy inference phase. We can incorporate the sampled noises into the MDP as part of the environment state.
Formally, for the original environment MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P_{0},P,R,\gamma\rangle$ we introduce MDP with noise-augmented states $\mathcal{M}_{noise}=\langle\mathcal{S}_{noise},\mathcal{A},P_{0},P_{noise},R,\gamma\rangle$. In $\mathcal{M}_{noise}$, each state $s_{noise}$ consists of a environment state $s\in\mathcal{S}$ and Gaussian noises $a^{K},z^{1},\cdots,z^{K}$. $\mathcal{M}_{noise}$ shares the same action space and reward function as the original MDP $\mathcal{M}$. In each decision-making step, the environment state transits to a new one, and the noises are all re-sampled.
Stochastic Noise-Conditioned Policy.
@@
Given a noise-augmented state $s_{noise}=(s,a^{K},z^{1:K})$, the deterministic inference phase of the denoising process can be represented as a Noise-Conditioned Policy $\pi^{NC}_{\theta}$ that generates the action $a^{0}$ using Eq. 7.
@@
This noise-conditoned policy is a deterministic policy and could not be directly trained with PPO since the policy loss (Eq. 2) relies on a stochastic policy. Therefore, we introduce an additional operation to transform this deterministic policy into a stochastic one.
For continuous actions we use a Gaussian policy centered at $f_{\theta}$:
$\pi^{\text{NC}}_{\theta}(\cdot\mid s,a^{K},z^{1:K})\;=\;\mathcal{N}\!\big(f_{\theta}(s,a^{K},z^{1:K}),\ \mathrm{diag}(\sigma_{\text{act}})^{2}\big),$ (8)
where $\sigma_{\text{act}}$ is a learnable vector of standard deviations. For discrete actions we treat $f_{\theta}(\cdot)$ as logits and apply a temperatureed softmax:
$\pi^{\text{NC}}_{\theta}(a^{0}=i\mid s,a^{K},z^{1:K})\propto\exp\!\big(f_{\theta}(s,a^{K},z^{1:K})_{i}/T\big).$ (9)
With this construction, PPO can optimize over the interactive action $a^{0}$ directly while the diffusion model deterministically generates the stochastic policy conditioned on pre-sampled noise.
5.2 Fine-tuning the Noise-Conditioned Policy with PPO
Under the formulation of NCDPO, at each environment timestep $t$ we first sample Gaussian noise $a^{K}_{t},z_{t}^{1:K}$ and then generate $a^{0}_{t}=f_{\theta}(s_{t},a^{K}_{t},z_{t}^{1:K})$ , and then sample the executed action from $\pi^{\text{NC}}_{\theta}$ as in Eq. (8) or (9). We store the tuple
$(s_{t},\ a^{K}_{t},\ z_{t}^{1:K},\ a^{0}_{t},\ r_{t},\ s_{t+1})$
in the experience buffer. During optimization we reuse the stored noises $a^{K}_{t},z_{t}^{1:K}$ to deterministically recompute $a^{0}_{t}$ and compute the PPO loss on those actions; gradients are backpropagated through the entire denoising computation $f_{\theta}$. This design lets PPO operate on environment actions while updating all denoising steps end-to-end.
We instantiate the PPO objective in the standard clipped form, optimizing the probability of interactive actions produced by the noise-conditioned policy:
@@ 5.2 Fine-tuning the Noise-Conditioned Policy with PPO
$\begin{split}L(\theta\mid\theta_{k})=\mathbb{E}_{\begin{subarray}{c}a_{t}^{0}\sim\pi^{\text{NC}}_{\theta_{k}}(\cdot\mid s_{t},a^{K}_{t},z_{t}^{1:K})\end{subarray}}\Bigg[\sum_{t}\min\bigg(&\frac{\pi^{\text{NC}}_{\theta}(a_{t}^{0}\mid s_{t},a^{K}_{t},z_{t}^{1:K})}{\pi^{\text{NC}}_{\theta_{k}}(a_{t}^{0}\mid s_{t},a^{K}_{t},z_{t}^{1:K})}\,A_{t},\\ &\text{clip}\!\bigg(\frac{\pi^{\text{NC}}_{\theta}(a_{t}^{0}\mid s_{t},a^{K}_{t},z_{t}^{1:K})}{\pi^{\text{NC}}_{\theta_{k}}(a_{t}^{0}\mid s_{t},a^{K}_{t},z_{t}^{1:K})},\,1-\epsilon,\,1+\epsilon\bigg)\,A_{t}\bigg)\Bigg],\end{split}$ (10)
Rollout / training pseudocode. For clarity, we provide a short pseudocode of the rollout and training loop used by NCDPO.
@@
As illustrated in Algo 16, in policy rollout process, each step begins by sampling a sequence of noises, which are then used by the Diffusion Policy to generate the corresponding action. These sampled noises are stored in the buffer. During training phase, the stored noises are reused to recompute the actions, enabling gradient backpropagation through the entire denoising process. This allows PPO to directly update all denoising steps of the diffusion policy.